OpenStudy (anonymous):

NEED HELP ASAP DUE TODAY! Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y A;The solution is (1, 5) B: The solution is (21, 0) C;There are an infinite number of solutions. D;There is no solution.

OpenStudy (anonymous):

im thinking it is B.

OpenStudy (anonymous):

okay can you help me with some more? :)

OpenStudy (anonymous):

sure...

OpenStudy (anonymous):

How many solutions does the equation have? 2(2x+5)=4(x+3) A:one solution B: infinite solutions c:no solution

OpenStudy (anonymous):

im thinking A.

OpenStudy (anonymous):

How many solutions does the equation have? a+5=1/5(15+5a) A:one solution B:infinite solutions C:no solution

OpenStudy (anonymous):

um A.

OpenStudy (anonymous):

im not good at math so don't blame me if they are wrong...

OpenStudy (anonymous):

lol its okay neither am I

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so was that the last question or no \[\large\color{blue}:\color{blue})\]